OpenStudy (baby_bear69):

What is the minimum value for g(x) = x^2 – 10x + 16?

OpenStudy (sweetburger):

-b/2a

OpenStudy (sweetburger):

thats going to be your vertex value

OpenStudy (baby_bear69):

So minimum value is 2?

OpenStudy (sweetburger):

Well its a point in the form (x,y). Also how did you get 2?

OpenStudy (baby_bear69):

I guessed XD

OpenStudy (sweetburger):

ax^2+bx+c is the general form. x=-b/2a is the lowest point in terms of x. Then plug you x value back into g(x) = x^2 – 10x + 16.